What is the final volume of 45.0 L of carbon dioxide gas at 21.2°C if the gas is heated to 91.9°C? Assume the pressure and the amount of gas are held constant.
1
Expert's answer
2020-02-14T04:36:36-0500

Solution.

"\frac{V1}{T1} = \frac{V2}{T2}"

"V2 = \frac{V1 \times T2}{T1}"

"V2 = \frac{45.0 \times 364.9}{294.2} = 55.81 \ L"

Answer:

V2 = 55.81 L